During an experiment, a scientist collects 0.051g of an unknown gas that has a pressure of 102 kPa, a temperature of 50.0 degrees Celsius and a volume of 66.4 mL. What is the molar mass of this unknown gas and IF it was discovered to be a gas of a nonmetal element, what element would it be (JUSTIFY YOUR ANSWER) *
1
Expert's answer
2020-06-29T07:52:02-0400

Let's write ideal gas law

"pV = \frac{mRT}{M} \\
M = \frac{mRT}{pV} = \frac{5.1\times10^{-5}kg *8.31\frac{J}{mol*K}*323K}{1.02\times10^5Pa*6.64\times10^{-5}L} \approx 2.02\times10^{-2}\frac{kg}{mol} = 20.2\frac{g}{mol}"

In the periodic table it's the molar mass of neon